The foci of a hyperbola
x^2/16 -(y-2)^2/9=1

Answer by akhilreddy90(1)

using the formulae for eccentricity we get e=4/5
then for standard hyperbola foci r =(ae,o) or (-ae,o)
but for this one, X=ae
i.e,x=4*4/5=4
thenY=o,but y=2
therfore,foci r (5,2),(-5,2)